Difference of an odd integer and three times the next consequetive odd integer is 24
Answer by ankor@dixie-net.com(12701)

Difference of an odd integer and three times the next consequetive odd integer is 24.
:
Let the odd integer = x
Then the next odd integer = (x+2)
:
3(x+2) - x = 24
:
3x + 6 - x = 24
3x - x = 24 - 6
2x = 18
x = 9 is the 1st odd integer.
:
:
Check: 3(11) - 9 = 24